the mass m of liquid varies directly as its volume V. If the mass of liquid in a cubical container 5 cm on a side is 375 g, find the mass of the liquid in a cubical container 3 cm on a side?
Answer by stanbon(75887)

the mass m of liquid varies directly as its volume V.
---
m = kV
----
Solve for "k" using "If the mass of liquid in a cubical container 5 cm on a side is 375 g".
---
375 = k*5^3
---
k = 3
---------
Equation:
m = 3V
---
Find the mass of the liquid in a cubical container 3 cm on a side?
m = 3*3^3 = 81 g
======================
Cheers,
Stan H.
----------------------